OpenStudy (anonymous):

:) help me pleasee.. A spherical storage tank has a diameter of 14 ft. How many cubic feet of water will it hold? ______ cubic ft.

jimthompson5910 (jim_thompson5910):

The radius is half the diameter Radius: r = d/2 = 14/2 = 7 So r = 7 Volume of a sphere: V = (4/3)*pi*r^3 V = (4/3)*pi*(7)^3 V = (4/3)*pi*343 V = 343*(4/3)*pi V = (1372/3)*pi V = 457.33333*pi V = 457.33333*(3.14159) V = 1436.7538 So it will hold approximately 1436.7538 cubic ft of water.

OpenStudy (anonymous):

i got that too but it says its wrong..

jimthompson5910 (jim_thompson5910):

Use more digits of pi to get a more accurate answer (if necessary)

jimthompson5910 (jim_thompson5910):

How many digits do they want you to round to?

OpenStudy (anonymous):

it just says.. in the end (use \[\Pi =\] 22/7

OpenStudy (anonymous):

so maybe put it in a fraction??

jimthompson5910 (jim_thompson5910):

oh thanks for the clarification

OpenStudy (anonymous):

and its just a box.. and you have to enter the digits so idk

jimthompson5910 (jim_thompson5910):

V = (4/3)*pi*r^3 V = (4/3)*pi*(7)^3 V = (4/3)*pi*343 V = 343*(4/3)*pi V = (1372/3)*pi V = (1372/3)*(22/7) V = (1372*22)/(3*7) V = 30184/21 V = 4312/3 V = 1437.333 So the volume is approximately 1437.333

OpenStudy (anonymous):

its not that either:l is there any way that you could put that as an improper fraction or so??

jimthompson5910 (jim_thompson5910):

Did you try the answer 4312/3 ?

OpenStudy (anonymous):

no.. ill try it hold on.. :)

OpenStudy (anonymous):

YESS ITS CORRECT :D <3 thank youuu thank youu thank youu:)

jimthompson5910 (jim_thompson5910):

You're welcome